Find all solutions if
0 ≤ x < 2π.
Use exact values only. (Enter your answers as a comma-separated list.)
cos 2x cos x − sin 2x sin x = square root of 2 over 2
x=?

Answer by lwsshak3(11628)

Find all solutions if
0 ≤ x < 2π.
Use exact values only. (Enter your answers as a comma-separated list.)
cos 2x cos x − sin 2x sin x = square root of 2 over 2
x=?
***
Identity: cos(x+y)=cosxcosy-sinxsiny
cos 2x cos x − sin 2x sin x = square root of 2 over 2
cos(2x+x)=√2/2
2x+x=π/4, 7π/4
3x=π/4, 7π/4
x=π/12, 7π/12